A salmon swims 24 km downstream from its birthplace to the ocean in three hours. The return trip upstream to spawn takes five hours. How fast does the salmon swim in its water?
Found 2 solutions by mananth, stanbon:
Answer by mananth(12270)

A salmon swims 24 km downstream from its birthplace to the ocean in three hours. The return trip upstream to spawn takes five hours. How fast does the salmon swim in its water?
salmon speed = x
cirrent speed =y
d/r=t
downstream
24/x+y=3
3x+3y=24
/3
x+y=8.................1
upstream
24/(x-y)=5
5x-5y=24..............2
multiply 1 by 5
5x+5y=40
add it to 2
10x=64
x= 6.4 salmon 's speed
current speed = 1.6 mph
...
m.ananth@hotmail.ca

Answer by stanbon(57424)

A salmon swims 24 km downstream from its birthplace to the ocean in three hours. The return trip upstream to spawn takes five hours. How fast does the salmon swim in its water?
---
rate downstream = 24/3 = 8 km/h
---
rate upstream = 24/5 = 4.8 km/h
----------
Equations:
s + c = 8 km/h
s - c = 4.8 km/h
----
Add and solve for "s":
2s = 12.8
---
salmon's swim rate = 6.4 km/h
================================
Cheers,
Stan H.
==========